OpenStudy (anonymous):

T + s/3 = R Solve for S. T'f ?!

OpenStudy (anonymous):

S=3(R-T)

OpenStudy (anonymous):

How ?!

OpenStudy (hba):

\[T+\frac{ s }{ 3 } =R \]

OpenStudy (hba):

\[\frac{ 3T+S }{ 3 }=R\]

OpenStudy (hba):

\[3T+S=3R\]

OpenStudy (hba):

\[S=3R-3T\]

OpenStudy (hba):

S=3(R-T)

OpenStudy (hba):

^ Answer @annie.kopliku